martai:

Dylan opened a credit card account with $700.00 of available credit. Now that he has made some purchases, Dylan's account only has $581.00 of available credit. What is the percentage decrease of the amount of available credit in Dylan's account? A. 20% B. 17% C. 83%

Vocaloid:

% decrease = (amount decreased)/(initial amount) * 100 %

martai:

wouldn't be b ?

Vocaloid:

Yes. The amount decreased from 700 to 581, so 700-581 = 119 (119 / 700)*100% = 17%

martai:

ok thx (: